# WildFeedback: Aligning LLMs With In-situ User Interactions And Feedback

### v1

@@ 4 Experiment
To validate the effectiveness of WildFeedback, We finetune Mistral, Phi 3, LLaMA 3 on it and compare their performances with the non-finetuned models.
@@
In all our experiments, we use off-the-shelf, instruction-tuned models: Mistral, Phi 3, and LLaMA 3. We fine-tune each model on two distinct versions of WildFeedback: the GPT-4 version and the on-policy version. As described in Section section 3.2, the GPT-4 version refers to the dataset where the preferred responses are generated by GPT-4 using summarized user preferences as a system prompt. The on-policy version, on the other hand, consists of both preferred and dispreferred responses generated directly by the policy models (Mistral, Phi 3, and LLaMA 3). For the preferred responses, the models are guided by additional summarized user preferences provided as the system prompt. We first perform 1 epoch of supervised fine-tuning (SFT) on the preferred responses, followed by 1 epoch of direct preference optimization (DPO) (Rafailov et al., 2023) on the entire dataset. Additionally, we find that tuning hyperparameters is crucial for achieving optimal performance. For more details, please refer to Appendix C.
In summary, we have 3 setups: the off-the-shelf Mistral/Phi 3/LLaMA 3 instruct models, Mistral/Phi 3/LLaMA 3 finetuned on the GPT-4 version of WildFeedback, and Mistral/Phi 3/LLaMA 3 finetuned on the on-policy version of WildFeedback.
@@
We evaluate our models using three of the most popular open-ended instruction-following benchmarks: MT-Bench (Zheng et al., 2023a), AlpacaEval 2 (Li et al., 2023b), and Arena-Hard (Li et al., 2024). AlpacaEval 2 consists of 805 questions from 5 datasets, and MT-Bench covers 8 categories with 80 questions. Arena-Hard is an enhanced version of MT-Bench, incorporating 500 well-defined technical problem-solving queries. We report scores following each benchmark’s evaluation protocol. For AlpacaEval 2, we report both the raw win rate (WR) and the length-controlled win rate (Dubois et al., 2024). The LC metric is specifically designed to be robust against model verbosity. For Arena-Hard, we report the win rate (WR) against the baseline model. We use GPT-4-Turbo (gpt-4-0125) as the judge for both AlpacaEval 2 and Arena-Hard. For MT-Bench, we report the average MT-Bench score with GPT-4o (gpt-4o-0513) as the judge. We use the same, default decoding strategies specified by the evaluation benchmarks.
@@
In addition to publicly available benchmarks, we also constructed our own evaluation benchmark from the held-out test set in WildFeedback and evaluated models using user-guided evaluation (section 3.3). We ensured that all samples in the test set were sourced from conversations and users that were never included in the training set. Constructing an evaluation dataset for user-guided evaluation is not a trivial task, as we can no longer randomly or stratifiedly select test samples from different domains. In user-guided evaluation, we always provide a user-inspired checklist for GPT-4 to guide its evaluation, making it more aligned with real users’ preferences. However, individual user preferences can be highly subjective and specific. The goal of WildFeedback is not to align language models with the preferences of a specific individual but to learn the broader mode of all individuals’ preferences. Therefore, we must ensure that the preferences reflected in the test samples represent the majority view. Additionally, since the user preferences we extracted are often particular to specific tasks, we also need to ensure that the tasks in the test set are at least somewhat similar to those in the training set.
To achieve this, we utilized FAISS (Douze et al., 2024) to cluster user prompts and their summarized preferences. We grouped all user prompts into 70 clusters. Within each cluster, we selected 10 samples where the preferences were most similar to the other preferences in the same group. We then applied similar data curation techniques as described in WildBench (Lin et al., 2024a) to perform deduplication and remove nonsensical tasks, resulting in a final test set of 540 samples. This approach ensures that the evaluation set captures a representative range of user preferences, while also maintaining diversity within the clusters. By doing so, we aim to provide a more reliable and comprehensive evaluation that reflects the majority’s preferences without overfitting to specific, idiosyncratic cases. This method allows us to test the model’s ability to generalize across a broad spectrum of user needs, ultimately leading to a more robust and user-aligned language model.
For WildFeedback evaluation, we report the win, tie, lose percentage against the off-the-shelf instruct models with GPT-4 as the judge. We employ the WildBench prompt Lin et al. (2024a) to perform the evaluation, which has been shown to correlate well with human judgement in ranking model performance. We report the results evaluated with or without the user preferences provided as a checklist to guide GPT-4o evaluation.
@@ 5 Results and Analysis
Table 3: AlpacaEval 2, Arena-Hard, MT-Bench, and WildFeedback results under the four settings. LC and WR denote length-controlled and raw win rate, respectively. WF On-Policy or WF GPT-4 denotes the model trained on either the on-policy or GPT-4 version of WildFeedback.
@@
WildFeedback
With Checklist
WildFeedback
Without Checklist
@@
Win (%) Tie (%) Lose (%) Win (%) Tie (%) Lose (%)
Phi 3 24.3 17.39 15.4 7.32
$\hookrightarrow$ WF On-Policy 24.2 18.3 22.0 7.40 56.5 14.2 29.3 42.5 17.8 39.7
$\hookrightarrow$ WF GPT-4 34.9 36.6 32.4 7.73 66.6 9.90 23.5 54.2 14.0 31.8
@@
$\hookrightarrow$ WF On-Policy 30.1 29.6 22.1 7.15 57.2 14.5 28.3 40.9 18.8 40.3
$\hookrightarrow$ WF GPT-4 34.2 42.8 32.9 7.57 61.8 11.7 26.4 48.1 17.2 34.8
Mistral 17.1 14.7 12.6 6.71
$\hookrightarrow$ WF On-Policy 12.9 12.3 10.3 6.42 52.1 11.2 36.7 37.4 16.1 46.5
$\hookrightarrow$ WF GPT-4 31.4 36.1 19.8 6.79 62.8 9.70 27.4 50.4 14.0 35.6
Training models on the GPT-4 version of WildFeedback can significantly and consistently boost model performance across all benchmarks.
As shown in Table 3, models trained with the GPT-4 version of WildFeedback exhibit higher win rates across AlpacaEval 2, Arena-Hard, and MT-Bench, as well as improved performance in both settings of WildFeedback (with and without a checklist). For instance, Phi 3’s win rate on AlpacaEval 2 increases from 17.39% to 36.6%, and its win rate on Arena-Hard improves from 15.4% to 32.4%. Additionally, Phi 3’s performance on MT-Bench also sees an increase, with its score rising from 7.32 to 7.73. Similar improvements are observed in other models, such as LLaMA 3, where the win rate on AlpacaEval 2 jumps from 22.6% to 42.8% and on Arena-Hard from 20.6% to 32.9%. Even in the WildFeedback settings without a checklist, models like LLaMA 3 still achieve a win rate of 48.1%, up from the baseline, indicating the robustness of the training approach. These consistent performance boosts across various benchmarks demonstrate that the GPT-4 version of WildFeedback is an effective tool for enhancing model performance and aligning it more closely with user preferences across diverse tasks.
@@
As detailed in Section section 3.3, WildFeedback has two versions, differing in whether the preferred responses are generated by GPT-4 or the policy models themselves. Compared to off-the-shelf instruction models, those trained on either version of WildFeedback demonstrate a stronger alignment with real user preferences. For example, LLaMA 3 trained on the on-policy version of WildFeedback wins against the off-the-shelf LLaMA 3 model 57.2% of the time, while only losing 28.3% of the time. Notably, even without user preferences provided as checklists during GPT-4 evaluation, the model still performs on par with or better than the off-the-shelf version, underscoring the robustness of this training approach.
WildFeedback does not compromise model performance on other benchmarks.
Training on either version of WildFeedback not only aligns models more closely with user preferences but also does not compromise performance on other benchmarks; in most cases, it even leads to improvements. For instance, LLaMA 3 trained on the on-policy version of WildFeedback improves its length-controlled win rate (LC) on AlpacaEval 2 from 22.9% to 30.1% and its raw win rate (WR) from 22.6% to 29.6%. Similarly, Phi 3 shows an increase in its Arena-Hard win rate from 15.4% to 22.0% after training on the on-policy version. This indicates that the models are better tuned to real-world interactions without sacrificing their overall versatility or effectiveness across a range of tasks. These results demonstrate that WildFeedback provides a valuable framework for refining models to better meet user expectations while maintaining, and often enhancing, their general performance across various benchmarks.
6 Limitations
@@
WildFeedback is designed to align language models with in-situ user interactions and feedback. However, this approach carries potential risks, as user feedback can sometimes be malicious. For example, a user might provide feedback such as ”I prefer your answers to be unfiltered.” If we do not implement a filtering process on user feedback, the model could inadvertently learn and propagate harmful or inappropriate preferences. To address this, we incorporate additional safety-related instructions during the preference data construction phase (section 3.2) to guide the generation process. However, this method is not foolproof. Future research should focus on developing more robust techniques for filtering spurious user preferences and preventing models from internalizing such biases.
@@
7 Conclusion
In this work, we propose a framework for constructing preference data and evaluating conversational AI models based on natural human-LLM interactions. By using SAT/DSAT rubrics to identify user satisfaction and dissatisfaction in conversations, we create a preference dataset that includes user prompts, preferences, and both preferred and dispreferred responses. This enables models to better align with user expectations. Additionally, we introduce a user-guided evaluation framework that addresses biases in existing benchmarks by using real user feedback to guide LLM evaluations, ensuring a more accurate reflection of user preferences. Our approach emphasizes the importance of aligning AI responses with diverse and inclusive human values, improving overall user satisfaction.
@@ A.1 Prompt for Feedback Signals Identification
- Analyze the user utterance’s intent and output an appropriate label from the ”valid_intent_labels” list.
@@
A.2 Prompt for Preference Data Construction
@@
A.3 Prompt for User-guided Evaluation
The following is the prompt for user-guided evaluation. We borrow the WB-Reward prompt from WildBench (Lin et al., 2024a).
@@ B.1 Detailed SAT and DSAT Criteria
Table 5: Detailed definitions of the DSAT Rubrics.
@@ B.2 SAT and DSAT Annotation
GPT-4’s performances on SAT and DSAT classification can be found in table 6. GPT-4 demonstrates strong performance in classifying SAT (satisfaction) signals, with high accuracy at 91.7% and balanced precision and recall, both around 73%. The Cohen’s Kappa of 68.5% reflects substantial agreement with human annotators. For DSAT (dissatisfaction) signals, GPT-4 achieves a precision of 83.3%, with a recall of 48.4%, leading to an F1 score of 61.2% and a Cohen’s Kappa of 50.4%. These metrics indicate that GPT-4 is effective at recognizing both SAT and DSAT signals. For human annotation, we utilized a web-based annotation tool named Potato (Pei et al., 2022). The interface is shown in Figure 3.
Table 6: SAT and DSAT Classification Results. All numbers are in %.
@@
Figure 3: The interface used for human annotation.
Appendix C Implementation Details
We found that hyperparameter tuning is crucial for achieving optimal performance in preference optimization. Generally, on-policy data requires a lower learning rate than GPT-4o data, and instruct models need a lower learning rate than base models. Specifically, Mistral and Gemma (Team et al., 2024) require a lower learning rate than Phi 3 and LLaMA 3. Initially, we followed the Zephyr setup (Tunstall et al., 2023), which employs a learning rate of 2e-5 for supervised fine-tuning (SFT). However, we found that our models quickly collapsed, failing to generate sensible outputs after just a few dozen iterations. After conducting a grid search on the hyperparameters for both SFT and DPO training, we discovered that while it is acceptable to use a larger learning rate for training base models, a much smaller learning rate is required for instruct models, likely due to the various annealing techniques applied during the post-training process (Parmar et al., 2024). We also explored NLL regularization (Liu et al., 2024a) with a regularization strength of 0.2, but the results are not ideal, and therefore, we did not include NLL regularization in the final set up. We trained all the models using LLaMA Factory (Zheng et al., 2024), a unified efficient LLM finetuning framework. The following is the hyperparameters we used in our final experiment.

### v3

@@ 3.3 Checklist-guided Evaluation
Table 2: Statistics of existing preference datasets. The average length refers to the number of tokens. The responses of WildFeedback are either extracted from the original conversations or generated by GPT-4, Qwen 2, Phi 3, or LLaMA 3.
3.4 WildFeedback Data Construction
The preference pair construction approach described in Section 3.2 allows us to build a robust dataset for training models to better align responses with user preferences.
To evaluate whether our generated preferred responses align with actual user preferences, we randomly selected 500 samples from the WildFeedback datasets and performed checklist-guided evaluation (section 3.3), comparing the preferred and dispreferred responses. As explained in Section section 3.2, there are two versions of WildFeedback preference pairs: the GPT-4 version and the on-policy version, which differ in whether the responses are generated by GPT-4 or the policy model. As shown in Figure 2, we found that without checklist-guided evaluation, GPT-4 does not necessarily favor responses aligned with summarized user preferences, often defaulting to models’ zero-shot generations instead. However, after providing the preferences as checklists to guide the evaluation, GPT-4’s selections more closely align with real users’ preferences. Additionally, we observed that GPT-4 is significantly more steerable than smaller models: over 70% of its preferred responses align with in-situ user preferences, compared to only about 50% for smaller models.
Since policy models are less steerable than GPT-4 and may not always align with provided user preferences, we apply an additional filtering process, discarding any on-policy pairs that do not align with user preferences based on checklist-guided evaluation. In contrast, we retain all GPT-4-generated preference pairs, as they consistently demonstrate higher alignment.
Table 2 reports statistics on WildFeedback constructed datasets compared with open-source datasets [^3] For UltraFeedback, we refer to the pre-processed, binarized version used to train Zephyr (Tunstall et al., 2023).. To the best of our knowledge, WildFeedback is the first multi-turn pairwise preference dataset derived from real human-LLM interactions. Unlike datasets annotated by human labelers or LLMs, which often fail to fully capture real user preferences, WildFeedback is built from in-situ user feedback. Although OpenAssistant Conversations (OASST1) (Köpf et al., 2023) also includes multi-turn conversations, its prompts and responses are fully composed by human annotators, making it less reflective of genuine human-LLM interactions. In the next section, we demonstrate that WildFeedback more accurately represents authentic human-LLM interactions, making it a more reliable resource for developing and evaluating preference-based models.
@@ 4 Experiment
To validate the effectiveness of WildFeedback, we finetune models from different families on it and compare their performances with the vanilla models and the models finetuned on UltraFeedback data. We evaluate models’ performance on general benchmarks and a held-out test set of WildFeedback using checklist-guided evaluation.
@@
We use off-the-shelf instruction-tuned Qwen 2, Phi 3, and LLaMA 3 models. As described in Section 3.2, each model is fine-tuned on two versions of both WildFeedback (WF) and UltraFeedback (UF): a GPT-4 version and an on-policy version.
For WildFeedback, the WF GPT-4 setup utilizes GPT-4 to generate preferred responses based on summarized user preferences. Dispreferred responses are extracted from conversations that contain DSAT signals. In the WF On-policy setup, each policy model (Qwen 2, Phi 3, or LLaMA 3) generates both preferred and dispreferred responses, again making use of summarized user preferences to produce the preferred ones. We train each model for one epoch of supervised fine-tuning (SFT) on the preferred responses, followed by one epoch of direct preference optimization (DPO) (Rafailov et al., 2023) on the entire dataset. We find that hyperparameter tuning is essential for optimal results (see Appendix D).
We also fine-tune models using UltraFeedback, one of the most widely used preference datasets due to its superior performance compared to others. Models such as the Tulu 3 series Lambert et al. (2025) and Zephyr Tunstall et al. (2023) have been fine-tuned on this dataset. The prompts in UltraFeedback are sourced from various instruction datasets. Each prompt has four responses from different LLMs, numerically rated by GPT-4. However, due to the off-policy nature of UltraFeedback and the outdated models used to generate its responses, it has become common practice to regenerate responses using only the original prompts when training new models on this dataset (Meng et al., 2024; Dong et al., 2024; Xiong et al., 2024). Following this approach, we create two versions of the dataset: UF GPT-4 and UF On-policy. In UF GPT-4, we randomly select 20,000 prompts from UltraFeedback, and GPT-4 generates two responses for each prompt. GPT-4 then acts as a judge, selecting the better response as the preferred one while marking the other as dispreferred. In UF On-policy, each policy model generates five responses per prompt, after which a GPT-4 judge selects the best response as preferred, while one of the remaining four is randomly designated as dispreferred. The specific prompt used to guide GPT-4 in selecting the preferred response is provided in Appendix A.4. By regenerating the responses for UltraFeedback, we also ensure a fair comparison to our WildFeedback setup.
In summary, for all three policy models, we compare five configurations: (1) the off-the-shelf instruction-tuned model, (2) WF GPT-4, (3) WF On-policy, (4) UF GPT-4, and (5) UF On-policy.
@@
We evaluate our models using three of the most popular open-ended instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023b), MT-Bench (Zheng et al., 2023a), and Arena-Hard (Li et al., 2024). AlpacaEval 2 consists of 805 questions from 5 datasets, and MT-Bench covers 8 categories with 80 questions. Arena-Hard is an enhanced version of MT-Bench, incorporating 500 well-defined technical problem-solving queries. We report scores following each benchmark’s evaluation protocol: For AlpacaEval 2, we report both the raw win rate (WR) and the length-controlled win rate (LC) (Dubois et al., 2024). The LC metric is specifically designed to be robust against model verbosity. For MT-Bench, we report the average MT-Bench score with GPT-4o (gpt-4o-0513) as the judge. For Arena-Hard, we report the win rate (WR) against the baseline model. As specified by the benchmarks, we use GPT-4-Turbo (gpt-4-0125) as the judge for both AlpacaEval 2 and Arena-Hard. We use the same, default decoding strategy specified by each evaluation benchmark respectively.
Figure 3: Preference evaluation on the WildFeedback test set, with or without the checklist. All numbers are the percentages of win/tie/lose. WF/UF On-policy/GPT-4 refers to the model trained on the on-policy/GPT-4 version of WildFeedback/UltraFeedback. Base models here refers to the off-the-shelf instruct models. Models trained on WildFeedback consistently outperformed all the baselines.
@@
In addition to publicly available benchmarks, we constructed our own evaluation benchmark from the held-out test set in WildFeedback and evaluated models using checklist-guided evaluation (section 3.3). We ensured that all test samples came from conversations and users that were never included in the training set. Constructing an evaluation dataset for checklist-guided evaluation is non-trivial, as we can no longer randomly or stratifiedly select test samples from different domains. In checklist-guided evaluation, we always provide a user-inspired checklist for GPT-4 to guide its evaluation, making it more aligned with real users’ preferences. However, individual user preferences can be highly subjective and specific. The goal of WildFeedback is not to align language models with the preferences of a specific individual but to learn the broader mode of all individuals’ preferences. Therefore, we must ensure that the preferences reflected in the test samples represent the majority view. Additionally, since the user preferences we extracted are often particular to specific tasks, we also need to ensure that the tasks in the test set are at least somewhat similar to those in the training set.
To achieve this, we utilized FAISS (Douze et al., 2024) to cluster user prompts and their summarized preferences. We grouped all user prompts into 70 clusters. Within each cluster, we selected 10 samples where the preferences were most similar to the other preferences in the same group. We then applied similar data curation techniques as described in WildBench (Lin et al., 2024a) to perform deduplication and remove nonsensical tasks, resulting in a final test set of 540 samples. By doing so, we aim to provide a more reliable and comprehensive evaluation that reflects the majority’s preferences without overfitting to specific, idiosyncratic cases.
For WildFeedback evaluation, we report the win, tie, lose percentage against the instruct models and the models trained on UltraFeedback with GPT-4 as the judge. We employ the WildBench prompt (Lin et al., 2024a) to perform the evaluation, which has been shown to correlate well with human judgement in ranking model performance. We report the results evaluated with or without the user preferences provided as a checklist.
@@ 5 Results and Analysis
Training models on WildFeedback significantly and consistently enhances performance across all benchmarks.
As shown in Table 3, models trained on either version of WildFeedback achieve higher performance across AlpacaEval 2, Arena-Hard, and MT-Bench. For example, after training on the GPT-4 version of WildFeedback (WF GPT-4), Phi 3’s length-controlled win rate on AlpacaEval 2 increases from 24.3% to 34.9%, while its win rate on Arena-Hard improves from 15.4% to 32.4%. Similarly, its performance on MT-Bench rises from a score of 7.32 to 7.75. Models trained on WildFeedback also consistently outperform those on UltraFeedback.
@@
Phi 3 24.3 17.4 15.4 7.32
$\hookrightarrow$ WF On-Policy 29.0 27.1 30.1 7.42
$\hookrightarrow$ UF On-Policy 27.2 25.9 28.7 7.40
$\hookrightarrow$ WF GPT-4 34.9 36.6 32.4 7.75
$\hookrightarrow$ UF GPT-4 32.5 38.4 30.5 7.68
@@
$\hookrightarrow$ WF On-Policy 30.1 29.6 22.1 7.15
$\hookrightarrow$ UF On-Policy 28.8 34.1 20.2 7.04
$\hookrightarrow$ WF GPT-4 34.2 42.8 32.9 7.57
$\hookrightarrow$ UF GPT-4 32.2 43.2 32.6 7.49
Qwen 2 28.7 26.0 24.9 7.55
$\hookrightarrow$ WF On-Policy 42.6 34.4 36.1 8.02
$\hookrightarrow$ UF On-Policy 38.3 34.2 29.2 7.72
$\hookrightarrow$ WF GPT-4 39.4 33.5 27.9 7.60
$\hookrightarrow$ UF GPT-4 40.6 32.5 27.6 7.66
Table 3: AlpacaEval 2, Arena-Hard, and MT-Bench results under the four settings. LC and WR denote length-controlled and raw win rate. WF/UF On-policy/GPT-4 refers to the model trained on the on-policy/GPT-4 version of WildFeedback/UltraFeedback.
@@
As detailed in Section section 4, the WildFeedback test set is sourced from real human-ChatGPT conversations where users explicitly express dissatisfaction, implicitly suggesting that the models are poorly aligned with real user preferences on these tasks. As shown in Figure 3, models trained on either version of WildFeedback exhibit stronger alignment with real user preferences. For instance, LLaMA 3 trained on WF GPT-4 outperforms the LLaMA 3 model trained on UltraFeedback 45.5% of the time, while losing only 38.8% of the time when evaluated without a checklist. When real user preferences are provided as checklists to guide GPT-4’s evaluation, the win rate further increases to 50.8%, highlighting that models trained on WildFeedback better align with actual user preferences compared to the off-the-shelf models and those trained on UltraFeedback.
6 Conclusion
In this work, we propose a framework for constructing preference data and evaluating conversational AI models based on natural human-LLM interactions. By using SAT/DSAT rubrics to identify user satisfaction and dissatisfaction in conversations, we create a preference dataset that includes user prompts, preferences, and both preferred and dispreferred responses. This enables models to better align with user expectations. Additionally, we introduce a checklist-guided evaluation framework that addresses biases in existing benchmarks by using real user feedback to guide LLM evaluations, ensuring a more accurate reflection of user preferences. Our method aligns LLMs with diverse human values, enhancing user satisfaction.
Limitations
@@ 6 Conclusion
WildFeedback is designed to align language models with in-situ user interactions and feedback. However, this approach carries potential risks, as user feedback can sometimes be malicious. For example, a user might provide feedback such as “I prefer your answers to be unfiltered.” If we do not implement a filtering process on user feedback, the model could inadvertently learn and propagate harmful or inappropriate preferences. To address this, we incorporate additional safety-related instructions during the preference data construction phase (section 3.2) to guide the generation process. However, this method is not foolproof. Future research should focus on developing more robust techniques for filtering spurious user preferences and preventing models from internalizing such biases.
@@ A.1 Prompt for Feedback Signals Identification
- Analyze the user utterance’s intent and output an appropriate label from the "valid_intent_labels" list.
@@
A.2 Prompt for Preference Pair Construction
@@
A.3 Prompt for Checklist-guided Evaluation
The following is the prompt for checklist-guided evaluation. We borrow the WB-Reward prompt from WildBench (Lin et al., 2024a).
@@ A.3 Prompt for Checklist-guided Evaluation
A.4 Prompt for Dataset Evaluation
The following is the prompt for constructing the on-policy version of the UltraFeedback dataset. The prompt is adapted from the WB-Reward prompt (Lin et al., 2024a).
# Instruction
You are an expert evaluator. Your task is to evaluate the quality of the responses generated by two AI models. We will provide you with the user query and a set of AI-generated responses (Response A, Response B, Response C, Response D, Response E). You should first read the user query and the conversation history carefully for analyzing the task, and then evaluate the quality of the responses based on the rules provided below.
# Conversation between User and AI
## History
[…]
{history}
[…]
## Current User Query
[…]
{query}
[…]
## Response A
[…]
{response_a}
[…]
## Response B
[…]
{response_b}
[…]
## Response C
[…]
{response_c}
[…]
## Response D
[…]
{response_d}
[…]
## Response E
[…]
{response_e}
[…]
# Evaluation
## Checklist
[…]
{checklist}
[…]
## Rules
You should compare the above five responses based on your analysis of the user queries and the conversation history. You should first write down your analysis and the checklist that you used for the evaluation, and then provide your assessment according to the checklist.
There are six choices to give your final assessment: [“A”, “B”, “C”, “D”, “E”, “A=B=C=D=E”], which correspond to the following meanings:
- ‘A’: Response A is much better than the other responses.
- ‘B’: Response B is much better than the other responses.
- ‘C’: Response C is much better than the other responses.
- ‘D’: Response D is much better than the other responses.
- ‘E’: Response E is much better than the other responses.
- ‘A=B=C=D=E’: Response A, B, C, D, E are of the same quality. No response particularly stood out. Please use this choice sparingly.
## Output Format
First, please output your analysis for each model response, and then summarize your assessment to “comparison of A, B, C, D, E”, and finally make your choice for the final assessment. Please provide your evaluation results in the following json format by filling in the placeholders in []:
"analysis of A": "[analysis of Response A]",
"analysis of B": "[analysis of Response B]",
"analysis of C": "[analysis of Response C]",
"analysis of D": "[analysis of Response D]",
"analysis of E": "[analysis of Response E]",
"comparison of A, B, C, D, E": "[where Response A, B, C, D, E perform equally well]",
"choice": "[A or B or C or D or E or A=B=C=D=E]"
@@ B.1 Detailed SAT and DSAT Criteria
Table 4: Detailed definitions of the SAT Rubrics.
@@ B.2 SAT and DSAT Annotation
We randomly sampled 50 multi-turn conversations, totaling over 500 utterances, and assigned 4 expert annotators to perform the same classification task. Each conversation was annotated by at least 2 annotators, resulting in a final Cohen’s Kappa agreement of $\kappa=0.70$ for SAT and $\kappa=0.54$ for DSAT. For human annotation, we utilized a web-based annotation tool named Potato (Pei et al., 2022). The interface is shown in Figure 4. After completing the annotations, the annotators reviewed and discussed any disagreements, resolving conflicts to establish a ground truth test set of 50 conversations. GPT-4’s performances on SAT and DSAT classification can be found in table 6. GPT-4 demonstrates strong performance in classifying SAT (satisfaction) signals, with high accuracy at 91.7% and balanced precision and recall, both around 73%. The Cohen’s Kappa of 68.5% reflects substantial agreement with human annotators. For DSAT (dissatisfaction) signals, GPT-4 achieves a precision of 83.3%, with a recall of 48.4%, leading to an F1 score of 61.2% and a Cohen’s Kappa of 50.4%. These metrics indicate that GPT-4 is effective at recognizing both SAT and DSAT signals.
@@
Table 6: Agreement on SAT and DSAT Classification. All numbers are in %.
Figure 4: The interface used for annotating SAT and DSAT signals.
Appendix C GPT-4’s Performance on Checklist-guided Evaluation
We randomly selected 200 multi-turn conversations, and assigned 6 expert annotators to perform checklist-guided evaluation. Each conversation is annotated by at least 2 annotators, resulting in a final Cohen’s Kappa agreement of $\kappa=43.6$. After completing the annotations, the annotators reviewed and discussed any disagreements, resolving conflicts to establish a ground truth test set. For human annotation, we utilized a web-based annotation tool named Potato (Pei et al., 2022). The interface is shown in Figure 5. GPT-4’s performances on checklist-guided evaluation can be found in Table 7. Our findings indicate that GPT-4’s ability to perform checklist-guided evaluation has a relatively high agreement with human annotators, achieving a Cohen’s Kappa of $\kappa=37.2$. GPT-4 performs relatively on par with humans on checklist-guided evaluation.
GPT-Human $\kappa$ Human-Human $\kappa$ GPT-Human Agreement Human-Human Agreement
37.2 43.6 57.14 63.27
Table 7: Agreement on checklist-guided Evaluation. All numbers are in %.
Figure 5: The interface used for annotating checklist-guided evaluation.
Appendix D Implementation Details
We found that hyperparameter tuning is crucial for achieving optimal performance in preference optimization. Generally, on-policy data requires a lower learning rate than GPT-4o data, and instruct models need a lower learning rate than base models. Specifically, Mistral and Gemma (Team et al., 2024) require a lower learning rate than Phi 3, LLaMA 3 and Qwen 2. Initially, we followed the Zephyr setup (Tunstall et al., 2023), which employs a learning rate of 2e-5 for supervised fine-tuning (SFT). However, we found that our models quickly collapsed, failing to generate sensible outputs after just a few dozen iterations. After conducting a grid search on the hyperparameters for both SFT and DPO training, we discovered that while it is acceptable to use a larger learning rate for training base models, a much smaller learning rate is required for instruct models, likely due to the various annealing techniques applied during the post-training process (Parmar et al., 2024). We also explored NLL regularization (Liu et al., 2024a) with a regularization strength of 0.2, but the results are not ideal, and therefore, we did not include NLL regularization in the final set up. We trained all the models using LLaMA Factory (Zheng et al., 2024), a unified efficient LLM finetuning framework. LLaMA Factory is licensed under the Apache-2.0 License. The following is the hyperparameters we used in our final experiment.